How do you write the ratio 8 to 3 1/6 as a fraction in lowest terms?
Answer by Alan3354(69443)

How do you write the ratio 8 to 3 1/6 as a fraction in lowest terms?
----------
3 1/6 = 19/6
8 = 48/6
------------
48/19